The average speed of a train in the onward journey is 25% more than that in the return journey. The train halts for one hour on reaching the destination. The total time taken for the complete to and from journey is 17 hours, covering a distance of 800 km. The speed of the train in the onward journey is:


[A] 45 km/hr


[B] 47.5 km/hr


[C] 52 km/hr


 


[D] 56.25 km/hr 

Grade:9

1 Answers

ganesh
10 years ago

the correct answer is (b) 47.5km/hr

             speed=distance/time

            speed=800/17+1

                     =800/18=47.5